OpenStudy (anonymous):

write the simplest polynomial function with the given zeros. 4i, 2, - 2

OpenStudy (anonymous):

get ready to multiply

OpenStudy (anonymous):

im ready!

OpenStudy (mathmale):

Let's start with an example. Suppose the given zeros are 1 and 2. Then the corresponding polynomial is y = (x-1)*(x-2). Which polynomial corresponds to the zeros 2, - 2?

OpenStudy (mathmale):

Note that there IS multiplication in the above example.

OpenStudy (anonymous):

(x-2)(x+2)

OpenStudy (anonymous):

= x^2 - 4

OpenStudy (anonymous):

yup that is one factor

OpenStudy (anonymous):

the other will be \((x+4i)(x-4i)=x^2+16\) multiply those two

OpenStudy (anonymous):



OpenStudy (anonymous):

you want me to check?

OpenStudy (anonymous):

yes pls

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

thank you guys! :-)